# Complementary Probes of Warped Extra Dimension: Colliders, Gravitational Waves and Primordial Black Holes from Phase Transitions

### v1

@@ Appendix A Thick-wall approximation
In this appendix, we derive the temperatures characterizing the phase transition. As we are interested in the large supercooling regime, we restrict ourselves to the $O(4)$ symmetric solution to the bounce equation in the thick-wall approximation. Within this approximation, the action of the bounce solution can be estimated as Megías et al. (2022)
@@
Fig. 9 shows in the plane $\{\rho,N\}$ the mismatch among the different approximations to evaluate $S_{n}$. For the solid blue lines, $S_{n}$ is computed numerically from Eq. (48); for the dotted black curve $S_{n}$ is approximated as in Eq. (51); and red lines correspond to the approximation $S_{n}=4\log(M_{P}/\rho)$ rather customary in the literature. As we can see, the typical values of $S_{n}$ are $100\lesssim S_{n}\lesssim 180$, and the approximation in Eq. (51) is small enough not to be visible. The disagreement from the naive estimation $S_{n}=4\log(M_{P}/\rho)$ is instead visible.
@@
$T_{n}$ $=\sqrt{\frac{3N}{S_{n}}}\left(\frac{\beta}{H_{\ast}}\right)^{1/4}\frac{\rho}{2\pi}\,,$ (52)
Moreover, by assuming an instantaneous reheating where all the energy of the potential is converted into radiation, and neglecting the nucleation temperature due to the supercooled regime, we compute the reheating temperature
@@
One can estimate $T_{p}$ as the temperature satisfying the condition $I(T_{p})=0.34$ which corresponds to $P\simeq 0.7$ Ellis et al. (2019); Lewicki et al. (2021). By plugging Eq. (43) into Eq. (56) in the vacuum dominated regime, the expression for $I(T)$ can be written as
@@ Appendix B Semi-analytical approximations
We have checked that such semianalytic approximations also work well for other values of $\rho$ and $N$ within the interval values considered in the present analysis. Equations. (65)-(68) and (64) are then the explicit expressions for the critical, nucleation, and reheating temperatures implemented in the main analysis of this work.

### v2

@@ Appendix A Thick-wall approximation
In this appendix, we derive the temperatures characterizing the phase transition. As we are interested in the large supercooling regime, we restrict ourselves to the $O(4)$ symmetric solution to the bounce equation in the thick-wall approximation. Within this approximation, the action of the bounce solution can be estimated as Megias:2021rgh
@@
Fig. 10 shows in the plane $\{\rho,N\}$ the mismatch among the different approximations to evaluate $S_{n}$. For the black dots, $S_{n}$ is computed numerically from Eq. (48); for the solid black curve $S_{n}$ is approximated as in Eq. (51); and red lines correspond to the approximation $S_{n}=4\log(M_{P}/\rho)$ rather customary in the literature. As we can see, the typical values of $S_{n}$ are $100\lesssim S_{n}\lesssim 180$, and the approximation in Eq. (51) is small enough not to be visible. The disagreement from the naive estimation $S_{n}=4\log(M_{P}/\rho)$ is instead visible.
@@
$T_{n}$ $=\sqrt{\frac{3N}{S_{n}}}\left(\frac{\beta}{H_{\ast}}\right)^{1/4}\frac{\rho}{2\pi}\,.$ (52)
Moreover, by assuming an instantaneous reheating where all the energy of the potential is converted into radiation  [^13] A naive condition for instantaneous reheating, that $\tau_{\chi}\lesssim t_{p}$, where $t_{p}$ is the age of the universe at percolation, leads to $\rho\gtrsim O(1)$ GeV, for $m_{\chi}=0.1\rho$. Smaller values of $\rho$ should be analyzed in more detail, which is beyond the scope of the present paper., and neglecting the nucleation temperature due to the supercooled regime, we compute the reheating temperature
@@
One can estimate $T_{p}$ as the temperature satisfying the condition $I(T_{p})=0.34$ which corresponds to $P\simeq 0.7$ Ellis:2018mja ; Lewicki:2021xku . By plugging Eq. (43) into Eq. (56) in the vacuum dominated regime, the expression for $I(T)$ can be written as
@@ Appendix B Semi-analytical approximations
We have checked that such semi-analytical approximations also work well for other values of $\rho$ and $N$ within the interval values considered in the present analysis. Equations. (65)-(68) and (64) are then the explicit expressions for the critical, nucleation, and reheating temperatures implemented in the main analysis of this work.